Xytex Products just paid a dividend of $2.00 per share, and the stock currently sells for...

  1. Xytex Products just paid a dividend of $2.00 per share, and the stock currently sells for $26. If the discount rate is 10 percent, what is the dividend growth rate?
Answer #1

Discount rate=(D1/Current price)+Growth rate

0.1=(2*(1+Growth Rate)/26+Growth Rate

0.1=(2+2*Growth Rate)/26+Growth Rate

0.1=0.076923076+0.076923076*Growth Rate+Growth Rate

Growth Rate=(0.1-0.076923076)/(0.076923076+1)

=2.14%(Approx).